### v2

# Classification of curl forces for all space dimensions

Abstract
We present a decomposition of classical potentials into a conservative (gradient) component and a non-conservative component. The latter generalizes the curl component of the force in the three-dimensional case. The force is transformed into a differential $1$-form, known as the work form. This work form is decomposed into an exact (gradient) component and an antiexact component, which in turn generalizes the curl part of the force. The antiexact component is subsequently decomposed using the Frobenius theorem. This local decomposition is a useful tool for identifying the specific components of classical potentials.
Keywords: curl force; positional force; circulatory force; generalized potential; nonconservative forces; generalized potential; antiexact forms; homotopy operator; geometric decomposition; Frobenius theorem;
Mathematical Subject Classification: 58A10, 53Z05, 51P05
1 Introduction
A well-known result in classical mechanics [24, 1, 12] is that a force with a vanishing curl is the gradient of a scalar function known as a potential. On a Riemannian manifold [2, 8, 20, 24, 15, 22, 1, 12, 23] $M$ with a metric tensor $g:TM\times TM\rightarrow\mathbb{R}$, a force can be represented by a differential $1$-form (the work form). In this framework, the result is equivalent to the statement that the work form is exact, meaning it is the exterior derivative of a 0-form (a scalar function), which is the potential. This property holds locally by the Poincaré lemma [10, 11, 21].
A force $F$ satisfying the condition $\nabla\times F\neq 0$ is known as a curl force, also referred to as a pseudo-gyroscopic or circulatory force. Such a force can still be decomposed into gradient-like components, although this decomposition is not unique. These forces are of physical significance [3, 5, 6] and can be quantized [7]. The classical motions under these forces can be classified [4]. They also imply also some topological properties of magnetic fields [16]. Therefore, understanding their structure is of paramount importance.
In [26] a decomposition of a curl force expressible by functions and their gradients was given, namely:
  • Proposition 2.3 [26]: In $\mathbb{R}^{2}$ a force $F:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ can be expressed as $F(x)=-V(x)\nabla U(x)$, where generalized potentials $U,V:\mathbb{R}^{2}\rightarrow\mathbb{R}$. Moreover, $\nabla V(x)=0\Leftrightarrow\nabla\times F=0$. More generally $V=f(U)\Leftrightarrow\nabla\times F=0$ for some function $f:\mathbb{R}\rightarrow\mathbb{R}$.
  • Proposition 2.4 [26]: In $\mathbb{R}^{3}$ a force $F:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ can be expressed as $F(x)=-V(x)\nabla U(x)-\nabla W(x)$, where generalized potentials $U,V,W:\mathbb{R}^{2}\rightarrow\mathbb{R}$, and $\nabla V\times\nabla U\neq 0$. Moreover, $\nabla V(x)=0\Leftrightarrow\nabla\times F=0$ and we can select $W=0$. We have $\nabla W=0\Leftrightarrow F\cdot\nabla\times F=0$.
These results stem from transforming [^1] The mappings $g:TM\rightarrow T^{*}M$ and $g^{-1}:T^{*}M\rightarrow TM$ are called musical isomorphisms [24, 1, 12]. a force $F\in TM$ into the work form $\omega=g(F,\_):TM\rightarrow\mathbb{R}$, i.e., $\omega\in T^{*}M$. Applying the Darboux theorem [10, 11] then determines the local form of $\omega$, which is subsequently transformed back into a vector field on the $TM$ bundle. Due to the use of the Darboux theorem, the resulting decomposition is local in nature and not unique. This approach is therefore distinct from the global Helmholtz-Hodge decomposition [20, 25], which is valid for compact manifolds. Moreover, these findings suggest that it is more natural to work with forces by expressing them as differential forms, particularly as the standard notion of a curl does not exist in higher-dimensional spaces.
In this paper, we propose a new, algorithmic approach to decompose any force into a potential component and a generalized curl component. The curl component is then non-uniquely decomposed into terms involving generalized potentials and their gradients using the Frobenius theorem. Unlike the method outlined in [26], our proposed decomposition does not require solving PDEs. This method provides a practical way of working with classical potentials in any number of dimensions. To address the difficulty of defining a curl outside of three-dimensional space, we replace it with the more general notion of an antiexact differential form. The proposed method also yields a more fine-grained decomposition than the one proposed in [26].
The paper is organized as follows. In the next section, we summarize the key results from the theory of exterior differential forms that are needed to construct the decomposition algorithm. We then present our main result and provide several examples that illustrate the algorithm’s application.
2 Preliminaries
In this section we summarize results from local behaviour of exterior differential forms following [10, 11, 21, 17, 18, 14]. They will be used for the construction of the force decomposition in the next section.
The main object of consideration is an $n$-dimensional smooth manifold $M$ equipped with a metric tensor $g:M\rightarrow\mathbb{R}$, since an inner product is defined in most physical problems. We will also use the exterior derivative $d:\Lambda^{k}(M)\rightarrow\Lambda^{k+1}(M)$.
On this manifold, we can associate a vector field $F\in TU$ with a $1$-form $\omega$ musical isomorphism $\omega=g(F,\_)$. Similarly, the gradient of a function $f:U\rightarrow\mathbb{R}$ defined via its differential $df$ such that $df=g(\nabla f,\_)$.
The curl operation is defined only in three-dimensional space. This is because the cross product, and consequently the curl, is specific to three-dimensional Euclidean space, where an oriented plane (a bivector) can be uniquely represented by a vector [9, 12]. This representation is difficult in curvilinear coordinates and not possible in higher-dimensional spaces. A good summary of this situation is the statement from [12]: ”To my knowledge, however, there is no reason for ever writing out the components of the vector curl A in curvilinear coordinates; if the expression curl A appears, it is a sure sign that the vector in question was not the contravariant A but rather the covariant vector (…)”. Therefore, to operate in spaces of any dimension, one must abandon the notion of the curl and switch to the more general notion of the exterior derivative acting on differential forms. We will see that the natural generalization of the curl component of a vector is an antiexact differential form, which we define below. We also must note that the Helmholtz decomposition can be constructed in vector calculus [13], where the construction of multivectors is needed.
2.1 Geometric decomposition
We restrict our analysis to a star-shaped subset of $M$ that lies within a single local chart, which can therefore be identified with a star-shaped subset $U\in\mathbb{R}^{n}$. The metric tensor $g:M\rightarrow\mathbb{R}$ can then be restricted to $U$ in a chart.
The star-shaped domain $U$ contains a central point $x_{0}\in U$ and admits a linear homotopy $F:[0,1]\times U\rightarrow U$, $F(t,x)=x_{0}+t(x-x_{0})$ which connects any point $x\in U$ with $x_{0}$. We define the linear homotopy operator as
$H\omega=\int_{0}^{1}i_{\mathcal{K}}\omega|_{F(t,x)}t^{k-1}dt,$ (1)
where $\mathcal{K}=(x-x_{0})^{i}\partial_{i}$, $\omega\in\Lambda^{k}(U)$ is a $k$-differential form on $U$, and $i:\Lambda^{k}(U)\rightarrow\Lambda^{k-1}(U)$ is the insertion operator $i_{\mathcal{K}}\omega=\omega(\mathcal{K},\ldots)$. One can use a different homotopy in the definition, however the linear one makes the operator nilpotent $H^{2}=0$. It has also properties $dHd=d$ and $HdH=H$. We have also homotopy invariance formula
$dH+Hd=I-s_{x_{0}}^{*},$ (2)
where $s_{x_{0}}^{*}$ is the pullback along the constant map $s_{x_{0}}:x_{0}\hookrightarrow U$. The map $s_{x_{0}}^{*}$ can be nonzero only on $\Lambda^{0}(U)$. Since we will be considering $1$-forms, the pullback $s_{x_{0}}^{*}$ vanishes, and from (2) it follows that $Hd+dH=I$ on $\Lambda^{k}(U)$ for $k>0$. Moreover, since $(dH)^{2}=dH$ and $(Hd)^{2}=Hd$, the operators $Hd$ and $dH$ are projectors, and they define a direct sum decomposition of $\Lambda(U)$ into the exact [^2] Since the domain is star-shaped, an exact from is also closed (i.e., an elements of $ker(d)$) by the Poincare lemma. vector space $\mathcal{E}^{k}(U)=ker(d)=im(dH)$, and the antiexact module over $\Lambda^{0}(U)=C^{\infty}(U)$ denoted by $\mathcal{A}^{k}(U)=ker(H)=im(Hd)=\{\omega\in\Lambda^{k}(U)|H\omega=0,\omega|_{x_{0}}=0\}$. For $f\in\Lambda^{0}(U)$ we also have decomposition [^3] If $f$ is a classical potential then the decomposition agrees with the global gauge freedom realized by the additive group $(\mathbb{R}=ker(d),+)$. $f(x)=\underbrace{(f(x)-f(x_{0})}_{\mathcal{A}}\oplus\underbrace{f(x_{0})}_{\mathcal{E}}$ as $f(x_{0})\in ker(d)$, $f(x)-f(x_{0})\in ker(H)$ and $(f(x)-f(x_{0}))|_{x_{0}}=0$. In summary, these statements yield the decomposition
$\Lambda^{k}(U)=\mathcal{E}^{k}(U)\oplus\mathcal{A}^{k}(U),~k\in\{0,\ldots,n\},$ (3)
which is called the geometric decomposition. It is local and associated with a specific choice of the homotopy center $x_{0}$. A different choice of $x_{0}$ induces a different decomposition; see [10]. The locality of the decomposition does not pose a problem when one is interested in a specific solution valid in a neighborhood of a point. Various applications of the geometric decomposition are presented in [10, 11, 21, 17, 18, 14].
The Poincare lemma can be proved using the geometric decomposition as $d\omega=0\Leftrightarrow\omega=dH\omega$, that is, if $\omega$ is closed, then it is also exact on a star-shaped subset.
We also remark that a similar construction exists on a Riemannian manifold for the codifferential operator $\delta:\Lambda^{k}(U)\rightarrow\Lambda^{k-1}(U)$, which is related to the exterior derivative via the Hodge star operator. A homotopy operator for the codifferential, known as the cohomotopy operator, also exists; see [18, 19].
2.2 Frobenius theorem
In this subsection, we summarize the main construction related to the Frobenius theorem, following [10]; see also [11, 22].The main idea is the notion of a (homogeneous) ideal $I$ of $\Lambda(U)$ that is a set closed with respect to the sum of forms and multiplication: $\alpha\wedge\beta\in I$ if $\beta\in I$ and for all $\alpha\in\Lambda(U)$. The ideal $I$ is a differential ideal, when $dI\subset I$, i.e., if the ideal is closed under exterior derivative. In many problems, the ideal is defined by a set of $1$-forms $\{\omega_{1},\ldots,\omega_{r}\}$ and is denoted by $I(\{\omega_{1},\ldots,\omega_{r}\})$. The exterior system of dimension $r$, $D_{r}$, is a set of $r$ linearly independent $1$-forms. It generates the ideal $I(D_{r})$. The system $D_{r}=\{\omega^{i}\}_{i=1}^{r}$ is completely integrable if and only if there exists $r$ independent functions $\{f^{i}(x)\}_{i=1}^{r}$ that each of the forms $\omega^{i}$ vanishes on each family of $r$-parameter (nr)(n-r)-dimensional surfaces $\{f^{i}(x)=c^{i}\}_{i=1}^{r}$ for $c^{i}\in\mathbb{R}$. This means, see [10], that $\omega^{i}=A^{i}_{j}df^{j}$ for some nonsingular $r\times r$ matrix $A(x)$ of functions.
We have
Theorem 1 (Frobenius, see [10] Theorem 4-3.2).
An exterior differential system $D_{r}$ defined by $r$ 11-forms $D_{r}=\{\omega_{i}\}_{i=1}^{r}$ is completely integrable if and only if the ideal $I(D_{r})$ is closed ($dI(D_{r})\subset I(D_{r})$), which is equivalent to
$d\omega^{i}=\Gamma^{i}_{j}\wedge\omega^{j},$ (4)
or
$\omega^{1}\wedge\ldots\wedge\omega^{r}\wedge d\omega^{i}=0,~i\in\{1,\ldots,r\},$ (5)
where $\Gamma$ is a $r\times r$ matrix of $1$-forms.
We will consider an ideal that is generated by a single $1$-form $\Omega$. We observe that in this case, the crucial role in the Frobenius theorem is played by the equation
$d\Omega=\Gamma\wedge\Omega+\Sigma,$ (6)
where $\Gamma\in\Lambda^{1}(U)$ and $\Sigma\in\Lambda^{2}(U)$. To this equation, one must adjoint its differential consequences
$d\Sigma=\Gamma\wedge\Sigma-\Theta\wedge\Omega,~d\Gamma=\Theta,~d\Theta=0,$ (7)
and treat them as a system of exterior differential equations. The form $\Sigma$ is called the torsion [^4] This is not the torsion from the Cartan moving frame method on Riemannian manifolds [22]. It describes only an obstruction to complete integrability of $\Omega$ with analogy to Riemannian curvature $2$-form for integrability of geodesics. It is standard equation of fundamental theories of physics - it is visible when it is rewritten as $d\Omega-\Gamma\wedge\Omega=\Sigma$ and identify the operator $d-\Gamma\wedge$\_ as a simplified version of exterior covariant derivative, see [19]. of the equation. The solution of this system is given by the following
Theorem 2 (see [10], Theorem 5-9.2).
Any system of exterior equations
$d\Omega=\Gamma\wedge\Omega+\Sigma,$ (8)
$d\Sigma=\Gamma\wedge\Sigma-\Theta\wedge\Omega,~d\Gamma=\Theta,~d\Theta=0,$ (9)
on $U$ is equivalent to the system
$d\Omega=\Gamma\wedge\Omega+\Sigma^{\prime},$ (10)
$d\Sigma^{\prime}=d\gamma\wedge\Sigma^{\prime},~\gamma=H(\Gamma),$ (11)
and has the general solution
$\Omega=e^{\gamma}(d\phi+\eta),~\Sigma^{\prime}=e^{\gamma}d\eta,$ (12)
with $\Sigma^{\prime}=\Sigma+H(d\Gamma)\wedge\Omega$,
$\phi=H(e^{-\gamma}\Omega),~\eta=H(e^{-\gamma}\Sigma^{\prime})=Hd(e^{-\gamma}\Omega),$ (13)
that are in $\mathcal{A}(U)$.
When torsion $\Sigma$ vanishes, we said that $\Omega$ is recursive with coefficient form $\Gamma$. In this case, the equations take the form
$\Omega=\Gamma\wedge\Omega,~\Theta\wedge\Omega=0,~d\Gamma=\Theta,$ (14)
with the solution
$\Omega=e^{\gamma}(d\phi+H(\theta\wedge d\phi)),$ (15)
where $\theta=Hd\Gamma$ with the constraint $d\theta\wedge\Omega=0$.
When, in addition to the recursive case, $\Gamma$ is exact, then $\Omega$ is called gradient recursive and the solution is
$\Omega=e^{\gamma}d\phi,~\phi=H(e^{-\gamma}\Omega).$ (16)
We note that in this case $e^{-\gamma}\Omega\in\mathcal{E}^{1}(U)$, i.e., $e^{-\gamma}$ is an integrating factor for $\Omega$.
We will call the formula for $\Omega$ the Frobenius decomposition for $\Omega$.
These results do not change if $\Omega$ is $k$-form for $k>1$; see [10] Section 5. One can also generalize these results to vector-valued differential forms; see [10, 19].
3 Force decomposition
We now combine geometric decomposition and the Frobenius theorem to decompose the work form into simpler components.
We begin by considering the work $1$-form $\omega$ corresponding to a force field $F:M\rightarrow TM$, defined as $\omega=g(F,\_)$. Since our method uses the homotopy operator, we restrict our analysis to a single chart, working within a star-shaped subset $U\subset\mathbb{R}^{n}$.
Our main result is the following theorem:
Theorem 3.
For a work force $\omega=g(F,\_)$ on a star-shaped subset $U\subset\mathbb{R}^{n}$ with a center $x_{0}\in U$, we have the decomposition
$\omega=df+\Omega,$ (17)
where $f=H\omega\in\Lambda^{0}(U)$ and $\Omega=Hd\omega\in\mathcal{A}^{1}(U)$.
If $\Omega\neq 0$ then we can further decompose the antiexact part $\Omega$ according to the structure of the equation $d\Omega=\Gamma\wedge\Omega+\Sigma$:
  • General case:
    $\Omega=e^{g}(dh+\eta),~\Sigma^{\prime}=e^{g}d\eta,$ (18)
    with $\Sigma^{\prime}=\Sigma+H(d\Gamma)\wedge\Omega$,
    $h=H(e^{-g}\Omega),~\eta=H(e^{-g}\Sigma^{\prime})=Hd(e^{-g}\Omega).$ (19)
  • Recursive case ($\Sigma=0$):
    $\Omega=e^{g}(dh+H(\theta\wedge dh)),$ (20)
    with the constraint $d\theta\wedge\Omega=0$.
  • Gradient recursive case ($\Sigma=0$, $\Gamma=dH\Gamma$):
    $\Omega=e^{g}dh,~h=H(e^{-g}\Omega).$ (21)
The functions $f,g,h$ can be called generalized potentials. The function $f$ corresponds to the standard potential when $\Omega=0$ when $g=0=h$.
Remark 1.
The form $\Omega\in\mathcal{A}^{1}(U)$ is a generalization of the curl part of the force.
Remark 2.
From historical reasons, the definition of potential used in physics is f-f. However, in this presentation, it is more natural to use this function without the minus sign.
Proof.
We use the notation and definitions from the previous section. The proof begins with the geometric decomposition of $\omega$:
$\omega=dH\omega+Hd\omega.$ (22)
We identify in the exact (’gradient’) part the potential $f=H\omega\in\Lambda^{0}(U)$. The second term, $\Omega=Hd\omega\in\Lambda^{1}(U)$, is the antiexact (’curl’) part as.
If $\Omega=0$, then $\omega=df$ is exact, and the force $F$ is a conservative gradient.
In the case where $\Omega\neq 0$, we have $d\Omega\neq 0$ since $\Omega$ is antiexact part (not in $ker(d)$). We can then apply the Frobenius decomposition to $\Omega$, which yields the result.
Transforming the results to the tangent space we obtain the following:
Corollary 1.
A vector $F:U\rightarrow TU$ on a star-shaped $U\subset\mathbb{R}^{n}$ with a center $x_{0}\in U$ can be decomposed as
$F=\nabla f+X,$ (23)
where, if $X\neq 0$ then
  • General case: $X=e^{g}\nabla h+Y$, where $Y=e^{g}g^{-1}(\eta,\_)$.
  • Recursive case: $X=\nabla h+e^{g}Y$, where $Y=g^{-1}(H(\theta\wedge dh),\_)$.
  • Gradient recursive case: $X=e^{g}\nabla h$.
Remark 3.
Note that for the general and recursive cases there is a vector $Y$ in the decomposition, which cannot be expressed as a gradient of a function.
The above corollary mixes covariant and contravariant objects, since the decomposition is performed on differential forms and then transformed back to the tangent space. The natural way to work with such a decomposition is at the level of differential forms.
From Theorem 3, we can infer the algorithm:
Algorithm 1.
Input: A force $F$, a star-shaped subset $U\in\mathbb{R}^{n}$, and a metric tensor $g$.
  1. 1.
    Compute the work form $\omega=g(F,\_)$.
  2. 2.
    Compute the exact part $dH\omega$ of $\omega$ and infer the potential $f=H\omega$.
  3. 3.
    Compute the antiexact part $\Omega=Hd\omega$.
  4. 4.
    If $\Omega\neq 0$:
    1. (a)
      Compute $d\Omega=dHd\Omega=d\omega$.
    2. (b)
      Select forms $\Gamma$ and $\Sigma$ that satisfy the equation $d\Omega=\Gamma\wedge\Omega+\Sigma$ (see Remark 4 on arbitrariness of this choice).
    3. (c)
      Compute $g=H\Gamma$, $h=H(e^{-g}\Omega)$, and $\eta=Hd(e^{-g}\Omega)$.
    4. (d)
      Construct the final decomposition: $\omega=df+e^{g}(dh+\eta)$.
  5. 5.
    else if $\Omega=0$:
    1. (a)
      Construct the decomposition $\omega=df$.
Remark 4.
The seemingly arbitrary selection of $\Gamma$ and $\Sigma$ for the case $\Omega\neq 0$ can be simplified in practice. As shown in Example 2 below, one can first select the factor $e^{-g}$, which fixes $\Gamma=dg$. Then, one can compute $h=H(e^{-g}\Omega)$ and $\eta=Hd(e^{-g}\Omega)$. In this approach, it is not necessary to compute $\Sigma$ explicitly.
We now provide an example of the application of this decomposition.
Example 1.
The simplest example is an exact work form. Consider $\mathbb{R}^{2}$ with the origin as its homotopy center and with the standard Euclidean metric. Take the work form
$\omega=xdx=d\left(\frac{1}{2}x^{2}\right).$ (24)
The potential is $f=\frac{1}{2}x^{2}$.
We can verify that this agrees with the geometric decomposition. We have
$dH\omega=d(\frac{1}{2}x^{2})=\omega,$ (25)
$Hd\omega=0=\omega=dH\omega,$ (26)
as required.
The decomposition $\omega=dH\omega+Hd\omega$ is therefore satisfied, with anriexact part being zero.
Example 2.
In this example, we consider an antiexact work form on $\mathbb{R}^{2}$ with the homotopy center at the origin and with the standard Euclidean metric. We select the form
$\omega=xdy-ydx.$ (27)
We find that $d\omega=2dx\wedge dy$, so the form is not exact. We apply the geometric decomposition
$dH\omega=d(\frac{1}{2}(xy-yx))=0,$ (28)
$\Omega=Hd\omega=\omega.$ (29)
Therefore, $\omega$ is purely antiexact.
We can now proceed with applying the (non-unique) Frobenius decomposition. We need to select $\Gamma$ and $\Sigma$ such that the equation $d\Omega=\Gamma\wedge\Omega+\Sigma$ is satisfied.A straightforward approach is to first select an factor $e^{-g}$, which fixes $\Gamma$, and then determine the corresponding $\Sigma$.
If we select $e^{-g}=x$, then $g=-\ln(x)$ and $\Gamma=dg=\frac{-1}{x}dx$. This gives $\Sigma=d\Omega-\Gamma\wedge\Omega=3dx\wedge dy$. We can then compute the potentials: $h=H(e^{-g}\Omega)=H(x^{2}dy-xydx)=0$. Since $d\Gamma=0$, we have $\Sigma^{\prime}=\Sigma$, and $\eta=H(e^{-g}\Sigma)=H(3xdx\wedge dy)=x^{2}dy-xydx=x\Omega$.
For an another choice, let $e^{-g}=x^{2}$. Analogous calculations yield $g=-\ln(x^{2})$, $\Gamma=\frac{-2}{x}dx$, $\Gamma\wedge\Omega=-2dx\wedge dy$, and $\Sigma=\Sigma^{\prime}=4dx\wedge dy$. The potentials in this case are $h=0$ and $\eta=x^{3}dy-x^{2}ydx$.
Comparing these two selections of the $e^{-g}$ term demonstrates that the Frobenius decomposition is not fixed; the corresponding terms $\Gamma$ and $\Sigma$ can be adjusted to find a case that is easier to calculate.
Example 3.
Consider a work form on $\mathbb{R}^{2}$, which we take to be star-shaped with the origin as its homotopy center. The work form is given by
$\omega=xdx+y^{2}dx.$ (30)
The first term xdxxdx is exact ($d(xdx)=0$) and is equal to $d(\frac{1}{2}x^{2})$. However, the second term is not exact, since $d(y^{2}dx)=2ydy\wedge dx\neq 0$.
First, we decompose $\omega$ into its exact and antiexact parts:
$dH\omega=d\left(\frac{1}{2}x^{2}+\frac{1}{3}xy^{2}\right)=xdx+\frac{2}{3}xydy+\frac{1}{3}y^{2}dx,$ (31)
$\Omega=Hd\omega=H(2ydy\wedge dx)=\frac{2}{3}(y^{2}dx-xydy)=\omega-dH\omega.$ (32)
Note that both terms of the original form $\omega$ contribute to exact part. contribute to the final exact part. This shows that even if a component of a form is not exact, it can still contribute to the exact part of the total form. The potential for exact term is
$f=\frac{1}{2}x^{2}+\frac{1}{3}xy^{2}.$ (33)
The next step is to apply Frobenius decomposition to $\Omega$. In this case, it is simpler to rearrange $\Omega$ directly:
$\Omega=\frac{2}{3}xy^{2}\left(\frac{dx}{x}-\frac{dy}{y}\right)=\frac{2}{3}xy^{2}d\left(\ln\left(\frac{x}{y}\right)\right).$ (34)
We therefore have $e^{g}=\frac{2}{3}xy^{2}$ and $h=\ln\left(\frac{x}{y}\right)$ and it is visible that it is a gradient recursive case.
4 Discussion
The exact part of the work form is associated with the conservative force, a classical result in mechanics. The antiexact part, as was noted in [19], can be thought of as the influence of an external agent (or a non-autonomous influence) from outside the system.
Its further decomposition using the Frobenius theorem extracts the integrable part, $e^{g}dh$, which can be interpreted in terms of inaccessibility, as in the Carathèodory interpretation of thermodynamics; see [26]. In the gradient recursive case, this is the only term comprising the antiexact part $\Omega$. It can be interpreted as a standard potential-driven force derived from $h$, but where the system’s coupling to this potential is modulated by a position-dependent scaling factor $e^{g}$. Physically, it could model phenomena like a charged particle moving in a magnetic field where the field strength varies in a specific way, or certain types of controlled robotic manipulations.
For the full and recursive case (but not for the gradient recursive case), the term $e^{g}\eta$ is also presents. It represents a more significant obstruction to integrability, as it cannot be written as the exterior derivative of a function multiplied by another function. This term represents the ”path-dependent core” of the force, which cannot be eliminated by a simple scaling factor. It can be present in driven systems where the external agent’s input is not derivable from any potential, scaled or otherwise. Note that this term is present in the full case (when torsion $\Sigma\neq 0$) or in the recursive case (when torsion vanishes, however when $\Gamma$ has antiexact part). It therefore measures the torsion and the antiexactness of $\Gamma$.
These terms open a new way to split the system’s dynamics among these different factors.
One can also rewrite the equation (6) for the antiexact part $\Omega$ of the work form as
$d^{\nabla}\Omega=\Sigma,$ (35)
where
$d^{\nabla}=d-\Gamma\wedge\_.$ (36)
The operator $d^{\nabla}$ bears a resemblance to the exterior covariant derivative on the trivial line bundle (or fiber set [19, 10]) $U\times\mathbb{R}\rightarrow U$. In general problem of parallel transport on a bundle, one provides the connection form $\Gamma$ and an external force $\Sigma$ to compute $\Omega$, as was presented in [19, 10]. However, for (6), we obtained the inverse problem: determine $\Gamma$ and $\Sigma$ for a fixed $\Omega$. This is a non-unique problem. Equation (35) also suggests that the antiexact force is a solution of a non-uniform parallel transport problem on a bundle with connection form $\Gamma$ and non-uniformity term $\Sigma$. As we infer in [19], the antiexact part $\Omega$ represents an external influence on the system, and equation (35) can be interpreted as an evolution of this external agent, seen through the force $\Omega$ with which it influence the system. Moreover, the curvature for this external system is given by [^5] In general the curvature of a connection form $\Gamma$ is given by $\Theta=d\Gamma+\Gamma\wedge\Gamma$. However, for a $\mathbb{R}$-valued form the term $\Gamma\wedge\Gamma$ vanishes. $\Theta=d\Gamma$.
5 Conclusions
A local decomposition for classical forces in a space of arbitrary dimension is proposed. This method employs the geometric decomposition on a star-shaped subset of a manifold, which distinguishes the exact part, associated with the gradient of a potential, from an antiexact part that generalizes the curl component of the force. The antiexact part is further resolved using the (non-unique) Frobenius decomposition into an integrable component and a remainder that cannot be represented as the exterior derivative of a function.
The proposed decomposition is constructive and can be applied to any local star-shaped subset of a manifold; an algorithm for its implementation has also been presented. A key advantage of this method is that it does not require solving partial differential equations (PDEs). Discussion of the possible meaning of each term is also provided.
Acknowledgments
The authors declare no conflict of interests.
No data is attached to the presented research.
The work was partially funded by the EU Horizon Europe MSCA grant No 101183077. The author would like to thanks Michael V. Berry and Pragya Shukla for enlightening discussion on the nature of curl forces.